muon piquet report week 9
Download
Skip this Video
Download Presentation
Muon Piquet Report week 9

Loading in 2 Seconds...

play fullscreen
1 / 5

Muon Piquet Report week 9 - PowerPoint PPT Presentation


  • 85 Views
  • Uploaded on

Muon Piquet Report week 9. Oleg Maev, 05.03.2012. Main activities. Oleg Maev, 05.03.2012. MUON starts to be running in Global after Winter shutdown. It was painful recovering after last (Monday) online intervention required restart of many PVSS projects, OPC-servers, etc.

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about 'Muon Piquet Report week 9' - fairly


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
muon piquet report week 9

Muon Piquet Report week 9

Oleg Maev, 05.03.2012

main activities

Main activities.

Oleg Maev, 05.03.2012

  • MUON starts to be running in Global after Winter shutdown.
  • It was painful recovering after last (Monday) online intervention required restart of many PVSS projects, OPC-servers, etc.
  • Something unusual is that the all global projects like MUSIDES, MUDAQA(C) were affected --- fixed by restart (many times for some of them).
  • Serious problem was only with OPC (CanOpen25+) server on both ODE PCs on A-side (muodea01w and muodea02w) – reboot of PC helped immediately for muodea02w and it seems cured the problem for muodea01w today. We had this kind of problem many times in the past on these PCs and very probably that there is a bug there.
main activities1

Main activities.

Oleg Maev, 05.03.2012

2. Threshold scan was done by Davide and recipes for widths 7 (~28ns) and 4 (~15ns) have been created. At present we are using the recipe for width 7. Most of tuned thresholds are commented (see comments on that next slide).

main activities2

Main activities.

Oleg Maev, 05.03.2012

On previous slide you can see some hot spots (noisy channels) randomly distributed in the system. I especially commented most of tuned threshold in AnomalousChannel.txt file. In fact it was almost empty at the moment of creating new recipes. I did it just to see if configuration of noise is really changed after interventions and I confirm that it is.

Apart of replacement of really noisy FEBs and removing another sources of noise (broken wires etc.) we have two main factors responsible for appearing of noisy channels:

1. Dark current – I’m sure that some of thresholds have been tuned due to this reason. Most of chambers with dark current are already cured and this source of noise is not present anymore.

2. Algorithm for threshold calculation sometimes gives strange values – It’s clearly seen again with new thresholds.

Practically, I need about one week to retune thresholds (It’s much faster with a new tool for searching noisy channels provided by Davide and Emiliano).

After that a new recipe for width 4(15ns) can be recreated and tested.

main activities3

Main activities.

Oleg Maev, 05.03.2012

3. Giovanni investigated on improvement with new version of TELL1 firmware.

4. Emiliano made the map of the LV marathon supply channels that control M1 FEBs. Now the state of the LV is visible.

https://lblogbook.cern.ch/MUON/3068

ad